OpenStudy (anonymous):

Franco bought 1.5 pounds of deli cheese for . If he went back for another 2 pounds of the same cheese, how much could Franco expect to pay ?

OpenStudy (anonymous):

U haven't put any figure, let's call it x.

OpenStudy (anonymous):

. +.25 .

OpenStudy (anonymous):

Let's ay it's a dollar, then

OpenStudy (anonymous):

its sapose to be$6.25 after for .

OpenStudy (anonymous):

\[6.25 + .25\times 6.25\]

OpenStudy (anonymous):

1.5/6.25 = 2/x 1.5x = 12.5 x = $8.33

OpenStudy (anonymous):

Divide it by 3 and get half a pound, multiply that by 4 to get 2 pounds...

OpenStudy (anonymous):

weird pricing system since if 1.5 pounds cost $6.25 then the price per pound is \[\frac{6.25}{1.5}=4\tfrac{1}{6}\] per pound. not likely

OpenStudy (anonymous):

In England it would cost a pound and fifty per pound...

OpenStudy (anonymous):

never seen that sign in any deli i have been to. @estudier really?

OpenStudy (anonymous):

Just playing with pounds...

OpenStudy (anonymous):

a pint's a pound...

OpenStudy (anonymous):

Europe is kilos nowadays, I forget when UK went over, 70 something, I think..

OpenStudy (anonymous):

So, u figure it out after all that,Lalakyshani?